Previous

Calendar

September

2007

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

26

JD238

27

JD239 (cool, overcast and rainy)
Site Visit (JB): Details

brought 4th N2 cylinder (CC199706) up to the trailer from N324

removed the TRAM trolley from its track for SO

got the Vaisala HMP35AC T/RH sensor (s/n S0630018) from the north canopy tower for the new aspiration unit

checked connections on the 21.5m C-SAT and into the FAST logger, couldn't see anything out of place...call Campbell

see details for newest cylinder pressure info

see details for HMP35AC 23x wiring scheme

28

JD240 (windy,cool,afternoon rain)
Site Visit (JB,JH,MP,NT): Details

POWER OUTAGE TO TOWER AND TRAILER FROM ~9:53-9:55 MDT

downloaded from RAD, PROF, MET, NCARTcs, SOIL, and Tcs

SpB Off

29

JD241
mailed broken AM25T back to Campbell (RMA#5257)

got new valve for Hewlett Packard N2 regulator #8507-0407, started new leak check (this valve was pro bono publico from AirGas)

30

JD242 (cool, heavy rain)
Site Visit (JB,NT): Details

Xcel Energy planned power outage at MRS from ~3:00-3:30 MDST

JB swapped 15 flasks in a heavy downpour with thunder and lightning

31

JD243
at 17:54 MDT: ndaq process dies, HYDRA data stop being collected.

JB Off

1

JD244
No HYDRA data.

2

JD245
No HYDRA data.

3

JD246
No HYDRA data.

JB back

4

JD247
No HYDRA data.

5

JD248
at ~16:15 MST: ndaq process re-started, HYDRA data start being collected again.

6

JD249 (cool, windy)
Site Visit (JB): Details

retired N2 cylinder CC17726 (~100 psi) and started N2 cylinder CC199706 (~2075 psi)

swapped 15 flasks

changed mag. perch. trap for the 6251

added salt and H2O to the grounding wells

Hydra flow rates looked fine

see details for Hydra cylinder pressures

brought N2 cylinders CC17726 and CC176258 (both ~100 psi) down to N324

rotated the shading band ~20 degrees counter clockwise, reset the PAR sensor into the plane of the shading band

SpB Returns

7

JD250

8

JD251

9

JD252

10

JD253
Foggy and misty from morning till afternoon (sonics not working well).

11

JD254

12

JD255 (clear, windy)
Site Visit (JB, MP, JH): Details

downloaded from cnr, prf, met, nctc, uc, ucb, soil, tc

@ 10:25 MDT set shading band to 9/12/07 solar noon sun angle (~36 degrees)

N2 9149741 (~500 psi) and N2 CC198686 (~300 psi) are also out at the tower

N2 cylinders CC193045, CC198701, CC125575 are the other new N2 cylinders at the trailer (new besides CC199706, which was just started at the tower on 9/6/07)

13

JD256
Site Visit (JB, SpB): Details , Photos (6)

swapped 15 flasks

took forest phenology photos

from 10-11:30 MST: Gast pump off at times; connected Nafion dryer using the spare pump in the trailer (old Nafion pump was dead).

at 11:35 MST: Connected UCOMO1/US3 to "B4" (previously B4 had a short tube that went about 1m away from Hydra box).

from 11:47-11:50 MST: ran cal gases through the sample cell. Also, turned down the flow rate for the ref cell (maybe too much!).

from 12:30-13:30 MST, reconnected repaired AM2T (sn 2283) to the soil logger.

at ~14:00 MST, made slight adjustment to the shading band...re-aligned in plane of band and tweaked slightly clockwise.

at ~14:30 MST, 2-d ati accidentally disconnected.

14

JD257

15

JD258

16

JD259

17

JD260
Yukon returned to CU motorpool.

18

JD261

19

JD262
Site Visit (JB, SL): Details

soil respiration

20

JD263 (warm/calm)
Site Visit (JB, SpB, SL, MP): Details , Photos (6)

swapped 15 flasks

took forest phenol photos

from 10:15-11:00 MST: did blow testing for most of the HYDRA inlets (see Details). Slightly increased the "ref" flow rate. Also, GPS'ed the locations of all inlets and TRAM towers.

at ~11:40 MST, 2-d ati (at 0.2m) re-connected.

21

JD264

22

JD265
Fall Equinox

23

JD266

24

JD267
Wet and rainy (morning).

25

JD268

26

JD269
Site Visit (JB, SpB, MP): Details , Photos (28)

replaced three ultra-torre fittings with swagelok unions on the profiler tubing going to the mini tower

started running ops45 (T5.met and RH5.met, and no high-rate ozone data)

discovered tubing for P_bar sensor was broken; fixed it using one of the severed sections

see details for tower cylinder info

jeep rental

27

JD270
Site Visit (JB, SLS): Details

pre-dawn water potential

jeep rental

28

JD271
Site Visit (JB, SL): Details

Stan took soil resp. measurements

swapped 15 flasks

replaced the cart wheels/tires

29

JD272

30

JD273
First significant snowfall (on Sat night).

1

JD274

2

JD275
More snow (Mon night) + strong winds.

3

JD276
Site Visit (JB, MP, AW, MW, JM, LL): Details

removed photo diode wands

4

JD277
Site Visit (JB, MP, SpB, NT, SL): Details , Photos (17)

reboot of russter2 ~10:00 MDT

swapped four flasks, done=2; switched robot program to semi-winter mode (seasonal flag=1, collects 4 flasks every week)

moved ATI 2-d sonic from 20cm to tower @ ~9m

from 13:40-14:10 MST: Finished blow-testing the HYDRA inlets

5

JD278

6

JD279